3. Five steps of ANOVA Suppose you are interested in how a persons marital status is related to his or her leisure activitieStep 1. Making Assumptions and Meeting Test Requirements Complete the following statements about the general assumptions requStep 2. Stating the Null Hypothesis For each population category (single, married, or divorced), let pus be the mean number oStep 3. Selecting the Sampling Distribution and Establishing the Critical Region The sampling distribution for ANOVA is the dStep 4. Computing the Test Statistic This step involves several computations, which you perform one at a time as follows: SSTStep 5. Making a Decision and Interpreting the Results Compare the test statistic with the critical value (enter the values y

PLEASE COMPLETE THE FOLLOWING FIVE STEPS OF ANOVA AND FILL IN THE BLANKS. THANKS IN ADVANCE
Answer #1

Step 1

Level of measurement is Nominal.

Populations are normally distributed

Population variances are same.

Step 2

Option B

H0: The average time spent watching TV or using the Internet is the same, regardless of martial status.

H1:

Option B

At least one of the population means is different

Step 3

Alpha = 0.01

N = 21

n = 7

Degress of Freedom:

dfBetween = a – 1 = 3-1 =2

dfWithin = N-a = 21-3 = 18

dfTotal = N-1 = 21-1 = 20

F Critical (dfBetween, dfWithin): (2,18) = 6.01

The degrees of freedom within is dfw = 18

The degrees of freedom between is dfb = 2

F Critical (dfBetween, dfWithin): (2,18) = 6.01

Step 4

SSBetwen = ∑(∑ai)2/n - T2/N = 40.10

SSWithin = ∑(Y)2 - ∑(∑ai)2/n = 26.86

SSTotal = SSBetwen ­­+ SSWithin = 66.95

MS = SS/df

MSBetween = SSBetween/dfBetween = 40.10/2 = 20.05

MSWithin = SSWithin/dfWithin = 26.86/18 = 1.49

F = MSeffect / MSerror = 20.05/1.49 = 13.44

Hence,

SST = 66.95; SSB = 40.10; SSW = 26.86;

MSBetween = 40.10/2 = 20.05

MSWithin = 26.86/18 = 1.49

F = 13.44

Step 5

Critical = 6.01

Obtained = 13.44

The test statistic falls outside the critical region, so your decision should be to reject the null hypothesis.

Option B

The average time spent watching TV or using the Internet differs significantly by martial status.